### v2

@@ 5.4 Register-Invite-Book System (RIB)
In the application of vaccine distribution, RIB has other noteworthy advantages. First, it is easy to prioritize specific groups of people. Those at higher risk, those prioritized for reasons of fairness, and those prioritized to prevent the spread of infection are likely to emerge for many diseases and vaccines, as discussed by Pathak2022reserve; Pathak2023Vaccine; Akbarpour2023prioritization. Prioritization criteria, such as infection risk or fairness, can matter even more than willingness to pay or exert effort. In RIB, policymakers can freely adjust the order in which invitations are sent to agents in the queue based on their observable attributes. Moreover, to distribute reservation slots among agents who belong to the same priority groups, SD implemented RIB saves the screening cost.
Second, RIB minimizes communication costs in the case of gradually supplied goods. Other mechanisms typically suffer from increased communication costs—for instance, when FCFS is used and new slots are added every midnight, participants have to repeatedly access the system every night until they successfully make a reservation, which imposes unnecessary effort on participants and increased server load. [^21] Such a system was indeed used in New York City during the COVID-19 pandemic, and people indeed checked the system right after midnight (verge_nyc_2021). Implementing SD as a direct mechanism would require creating a new lottery notification each time vaccines are supplied and having participants submit their preferences for new dates, which incurs high unmodeled costs. In contrast, with RIB, participants only need to access the system twice, for registration and booking. This property greatly reduces communication costs, even compared with FCFS.
@@ 6 Concluding Remarks
Second, we applied the automated mechanism design method established by dutting2019optimal to numerically explore efficient money-burning mechanisms. The results demonstrated that several strong assumptions imposed in the stylized environment—such as a very large variety of goods, a continuum of agents and objects, and uniform capacities—are not essential for the trend that SD tends to be efficient in multi-object allocations. This suggests that SD can perform well in practical settings. Given the various excellent properties of SD discussed in prior studies, our findings position SD as a promising option.
@@ A.2 Asymmetric Capacity
The efficient mechanism is neither SD nor VCG. The mechanism offers two options: Receiving object $1$ with high probability or receiving object $2$ with low probability, without requiring payments. Following the terminology of goldner2024simple, we call this mechanism the random-favorites mechanism (RF).222222goldner2024simple study an (ex ante) symmetric environment, and therefore, each agent $i$ claims $k\in\operatorname*{arg\,max}v_{k}^{i}$ as their favorite object in equilibrium. This property does not necessarily hold with asymmetric capacity because the probability of being assigned depends on which object to claim. By requiring which object to claim before drawing a lottery, this mechanism partially screens for agents’ preference for their favorite object relative to the other one—only agents who strongly prefer object $2$ over object $1$ desire to claim object $2$, given that object $1$ is much easier to obtain. By contrast, under SD, claiming object $2$ when it is available is risk-free, and therefore, all agents who prefer object $2$ over object $1$ obtain object $2$ whenever possible. Consequently, RF outperforms SD with asymmetric capacities.
While RF outperforms SD in the residual surplus, it comes with several drawbacks. While RF can be implemented as a strategy-proof mechanism in a continuous market, to determine the allocation probability of each option, we need information about the value distribution, which is often considered challenging in the market design literature (wilson1985game). By contrast, SD is detail-free. Furthermore, RF cannot satisfy strategy-proofness in a finite market, even if the market is large. [^23] Several studies on matching theory (e.g., abdulkadirouglu2011resolving) have demonstrated that non-strategy-proof mechanisms may outperform strategy-proof mechanisms in equilibrium because their equilibrium strategies can reflect agents’ preference intensity. Note also that RF may not satisfy even a weaker notion of truthfulness, such as Bayesian incentive compatibility (BIC). While goldner2024simple proves that RF satisfies BIC when each object $k$ has a unit capacity and each value $v_{k}^{i}$ is drawn i.i.d., these assumptions are rarely satisfied practically. This is because the object an agent should claim depends on which objects other agents are claiming. In addition, RF cannot satisfy the practically desirable properties that FCFS and the sequential-form SD meet (discussed in Section 5).
Figure A.2: Percentage difference of the residual surplus achieved by RF and SD $(RS(\mathcal{M}_{RF})-RS(\mathcal{M}_{SD}))/RS(\mathcal{M}_{RF})$.
@@ A.4 Performance of SD, VCG, and the (numerically) efficient mechanisms
For the single-good and continuous market cases, the efficient mechanism is derived analytically using the results from Hartline2008 and Section 3. In contrast, for the $c=1$ and $c=5$ cases, the mechanism is obtained numerically using RegretNet.
@@ B.1 Proof of Theorem 1
Since the number of pairs is finite, there exists a pair $(k^{-},k^{+})$ such that $F(V_{k^{-},k^{+}})>0$. Fix such a pair and write $V^{**}=V_{k^{-},k^{+}}$. For every $\mathbf{v}\in V^{*}$, $k^{-}(\mathbf{v})=k^{-}$ and $k^{+}(\mathbf{v})=k^{+}$.
@@ B.2 Proof of Theorem 4
Proof.
To this end, we will prove
for all $v\in[t,\bar{v})$.
$\frac{a^{K}(1-b^{K})}{1-a^{K}}-\frac{a^{K+1}(1-b^{K+1})}{1-a^{K+1}}\geq 0.$ (55)
This inequality is equivalent to
$(1-b^{K})(1-a^{K+1})-a(1-b^{K+1})(1-a^{K})\geq 0,$ (56)
or
Let $A=1+a+\dots+a^{K-1}$, and $B=1+b+\dots+b^{K-1}$. Then, (57) becomes
$B(1+aA)-a(1+bB)A\geq 0,$ (58)
or equivalently,
$(B-aA)+aAB(1-b)\geq 0.$ (59)
B.3 Proof of Theorem 2
We apply the construction of the ironed virtual value established in Section 6 of myerson1981optimal to a distribution function $G_{K}$.
@@ B.4 Proof of Theorem 5
We show that for all $v\in[0,\bar{v}]$, there exists $K_{0}$ such that for all $K>K_{0}$, $r^{\prime}(v;G_{K})>0$. The right-hand side of (71) is bounded above by $((g(v))^{2}-g^{\prime}(v)G(v))$. Since we assume that $G$ has full support, i.e., $g(v)>0$ for all $v\in(0,\bar{v})$, there exists $K_{0}$ such that
@@ B.6 Proof of Theorem 7
This lemma is a basic step in deriving revenue-maximizing mechanisms and has been well-known. We include it here for completeness. The proof follows the approach outlined in Fu2017myerson.
@@ B.7 Proof of Proposition 4
We first prove the following lemma.
@@
When the reduced value distribution $\hat{G}_{K}$ has an IDHR, the ironed virtual value $\bar{\vartheta}$ is given by
$\bar{\vartheta}(w;\hat{G}_{K})=\begin{cases}\vartheta(w^{**};\hat{G}_{K})&\text{ for }w\in[0,w^{**}]\\ \vartheta(w;\hat{G}_{K})&\text{ for }w\in(w^{**},+\infty)\end{cases}$ (102)
where $w^{**}$ is the solution of
$\frac{\displaystyle\int_{0}^{w}\vartheta(w^{\prime};\hat{G}_{K})\hat{g}_{K}(w^{\prime})dw^{\prime}}{\hat{G}_{K}(w)}=\vartheta(w;\hat{G}_{K}).$ (103)
@@
Since $\hat{G}_{K}$ has an IDHR, $H$ is concave for $0,\hat{G}_{K}(w^{*}))$ and convex for $(\hat{G}_{K}(w^{*}),+\infty)$. Accordingly, there exists a unique line passing through the origin and being tangent to $H$. Let the $q$-coordinate of the point of tangency be $q^{**}$. Then, the convex hull $I$ of $H$ is given by
$I(q)=\begin{cases}\dfrac{H(q^{**})}{q^{**}}\cdot q&\text{ for }q<q^{**}\\ H(q)&\text{ for }q\geq q^{**},\end{cases}$ (104)
implying that the ironed virtual value is given by (102), where $w^{**}=\hat{G}_{K}^{-1}(q^{**})$.
@@
$\frac{\int_{0}^{q}\vartheta(\hat{G}_{K}^{-1}(r);\hat{G}_{K})dr}{q}=\vartheta(\hat{G}_{K}^{-1}(q)).$ (105)
@@
By Theorem 2.8 of Hartline2008, an efficient mechanism maximizes the value of $\hat{x}(w)$ in regions where $\bar{\vartheta}(w;\hat{G}_{K})$ is larger while fulfilling the resource constraint and the unit demand condition. The conclusion is immediate from this. ∎
@@ B.8 Proof of Theorem 8
[…]
For the same reasons used in the proof of Lemma 3, $w^{{\dagger}}$ must solve $\Theta(w;\hat{G}_{K})=0$, and $w^{**}$ must solve $\Theta(w;\Phi_{\alpha})=0$. Since $\Phi_{\alpha}$ has an IDHR, $\Theta(\cdot;\Phi_{\alpha})$ is decreasing in the neighborhood of $w^{**}$. Furthermore, for any $\delta^{\prime\prime}>0$, by taking sufficiently large $K$, we can make
$|\Theta(w;\hat{G}_{K})-\Theta(w;\Phi_{\alpha})|<\delta^{\prime\prime}.$ (107)
Accordingly, for any $\epsilon>0$, for sufficiently large $K$, $w^{{\dagger}}\in(w^{**}-\epsilon,w^{**}+\epsilon)$. Together with the fact that $\hat{G}_{K}(w)\to\Phi_{\alpha}(w)$ uniformly in $\underline{c},\bar{c}]$, we have the desired conclusion. ∎

### v3

@@ 5.4 Register-Invite-Book System (RIB)
In the application of vaccine distribution, RIB has other noteworthy advantages. First, it is easy to prioritize specific groups of people. Those at higher risk, those prioritized for reasons of fairness, and those prioritized to prevent the spread of infection are likely to emerge for many diseases and vaccines, as discussed by Pathak et al. (2022, 2024); Akbarpour et al. (2023a). Prioritization criteria, such as infection risk or fairness, can matter even more than willingness to pay or exert effort. In RIB, policymakers can freely adjust the order in which invitations are sent to agents in the queue based on their observable attributes. Moreover, to distribute reservation slots among agents who belong to the same priority groups, SD implemented RIB saves the screening cost.
Second, RIB minimizes communication costs in the case of gradually supplied goods. Other mechanisms typically suffer from increased communication costs—for instance, when FCFS is used and new slots are added every midnight, participants have to repeatedly access the system every night until they successfully make a reservation, which imposes unnecessary effort on participants and increased server load. [^21] Such a system was indeed used in New York City during the COVID-19 pandemic, and people indeed checked the system right after midnight (Krasnoff, 2021). Implementing SD as a direct mechanism would require creating a new lottery notification each time vaccines are supplied and having participants submit their preferences for new dates, which incurs high unmodeled costs. In contrast, with RIB, participants only need to access the system twice, for registration and booking. This property greatly reduces communication costs, even compared with FCFS.
@@ 6 Concluding Remarks
Second, we applied the automated mechanism design method established by Dütting et al. (2019) to numerically explore efficient money-burning mechanisms. The results demonstrated that several strong assumptions imposed in the stylized environment—such as a very large variety of goods, a continuum of agents and objects, and uniform capacities—are not essential for the trend that SD tends to be efficient in multi-object allocations. This suggests that SD can perform well in practical settings. Given the various excellent properties of SD discussed in prior studies, our findings position SD as a promising option.
@@ A.2 Asymmetric Capacity
The efficient mechanism is neither SD nor VCG. The mechanism offers two options: Receiving object $1$ with high probability or receiving object $2$ with low probability, without requiring payments. Following the terminology of Goldner and Lundy (2024), we call this mechanism the random-favorites mechanism (RF). [^22] Goldner and Lundy (2024) study an (ex ante) symmetric environment, and therefore, each agent $i$ claims $k\in\operatorname*{arg\,max}v_{k}^{i}$ as their favorite object in equilibrium. This property does not necessarily hold with asymmetric capacity because the probability of being assigned depends on which object to claim. By requiring which object to claim before drawing a lottery, this mechanism partially screens for agents’ preference for their favorite object relative to the other one—only agents who strongly prefer object $2$ over object $1$ desire to claim object $2$, given that object $1$ is much easier to obtain. By contrast, under SD, claiming object $2$ when it is available is risk-free, and therefore, all agents who prefer object $2$ over object $1$ obtain object $2$ whenever possible. Consequently, RF outperforms SD with asymmetric capacities.
While RF outperforms SD in the residual surplus, it comes with several drawbacks. While RF can be implemented as a strategy-proof mechanism in a continuous market, to determine the allocation probability of each option, we need information about the value distribution, which is often considered challenging in the market design literature (Wilson, 1985). By contrast, SD is detail-free. Furthermore, RF cannot satisfy strategy-proofness in a finite market, even if the market is large. [^23] Several studies on matching theory (e.g., Abdulkadiroğlu et al., 2011) have demonstrated that non-strategy-proof mechanisms may outperform strategy-proof mechanisms in equilibrium because their equilibrium strategies can reflect agents’ preference intensity. Note also that RF may not satisfy even a weaker notion of truthfulness, such as Bayesian incentive compatibility (BIC). While Goldner and Lundy (2024) proves that RF satisfies BIC when each object $k$ has a unit capacity and each value $v_{k}^{i}$ is drawn i.i.d., these assumptions are rarely satisfied practically. This is because the object an agent should claim depends on which objects other agents are claiming. In addition, RF cannot satisfy the practically desirable properties that FCFS and the sequential-form SD meet (discussed in Section 5).
Figure A.2: Percentage difference of the residual surplus achieved by RF and SD $(RS(\mathcal{M}_{RF})-RS(\mathcal{M}_{SD}))/RS(\mathcal{M}_{RF})$ for which the RF mechanism admits an interior solution.
@@ A.4 Performance of SD, VCG, and the (numerically) efficient mechanisms
For the single-good and continuous market cases, the efficient mechanism is derived analytically using the results from Hartline and Roughgarden (2008) and Section 3. In contrast, for the $c=1$ and $c=5$ cases, the mechanism is obtained numerically using RegretNet.
@@ B.1 Proof of Theorem 1
Since the number of pairs is finite, there exists a pair $(k^{-},k^{+})$ such that $F(V_{k^{-},k^{+}})>0$. Fix such a pair and write $V^{**}=V_{k^{-},k^{+}}$. For every $\mathbf{v}\in V^{**}$, $k^{-}(\mathbf{v})=k^{-}$ and $k^{+}(\mathbf{v})=k^{+}$.
@@ B.2 Proof of Theorem 2
We apply the construction of the ironed virtual value established in Section 6 of Myerson (1981) to a distribution function $G_{K}$.
@@
B.3 Proof of Theorem 4
Proof.
To this end, we will prove
for all $v\in[t,\bar{v})$.
$\frac{a^{K}(1-b^{K})}{1-a^{K}}-\frac{a^{K+1}(1-b^{K+1})}{1-a^{K+1}}\geq 0.$ (64)
This inequality is equivalent to
$(1-b^{K})(1-a^{K+1})-a(1-b^{K+1})(1-a^{K})\geq 0,$ (65)
or
Let $A=1+a+\dots+a^{K-1}$, and $B=1+b+\dots+b^{K-1}$. Then, (66) becomes
$B(1+aA)-a(1+bB)A\geq 0,$ (67)
or equivalently,
$(B-aA)+aAB(1-b)\geq 0.$ (68)
@@ B.4 Proof of Theorem 5
We show that for all $v\in(0,\bar{v})$, there exists $K_{0}$ such that for all $K>K_{0}$, $r^{\prime}(v;G_{K})>0$. The right-hand side of (71) is bounded above by $((g(v))^{2}-g^{\prime}(v)G(v))$. Since we assume that $G$ has full support, i.e., $g(v)>0$ for all $v\in(0,\bar{v})$, there exists $K_{0}$ such that
@@ B.6 Proof of Theorem 7
This lemma is a basic step in deriving revenue-maximizing mechanisms and has been well-known. We include it here for completeness. The proof follows the approach outlined in Fu (2017).
@@ B.7 Proof of Proposition 4
If $\hat{G}_{K}$ is NBUE, then by Theorem 2, a no-screening mechanism is efficient. In this case, we may set $w^{**}=+\infty$ and take $\hat{x}(w)\equiv\bar{m}$.
In the following, we consider the case in which $\hat{G}_{K}$ is not NBUE. We first prove the following lemma.
@@
Suppose that the reduced value distribution $\hat{G}_{K}$ has an IDHR and is not NBUE. Then, the ironed virtual value $\bar{\vartheta}$ is given by
$\bar{\vartheta}(w;\hat{G}_{K})=\begin{cases}\vartheta(w^{**};\hat{G}_{K})&\text{ for }w\in(-\infty,w^{**}]\\ \vartheta(w;\hat{G}_{K})&\text{ for }w\in(w^{**},+\infty)\end{cases}$ (102)
where $w^{**}$ is the solution of the following equation:
$\frac{\displaystyle\int_{-\infty}^{w}\vartheta(w^{\prime};\hat{G}_{K})\hat{g}_{K}(w^{\prime})dw^{\prime}}{\hat{G}_{K}(w)}=\vartheta(w;\hat{G}_{K}).$ (103)
@@
Since $\hat{G}_{K}$ has an IDHR, $H$ is concave for $(0,\hat{G}_{K}(w^{*}))$ and convex for $(\hat{G}_{K}(w^{*}),1)$. Since $\hat{G}_{K}$ is not NBUE, the convex hull $I$ cannot coincide with the line segment connecting $(0,0)$ and $(1,H(1))$, and therefore the supporting line from the origin touches $H$ at an interior point $q^{**}\in(0,1)$; thus, we have the following:
$I(q)=\begin{cases}\dfrac{H(q^{**})}{q^{**}}\cdot q&\text{ for }q<q^{**},\\ H(q)&\text{ for }q\geq q^{**}.\end{cases}$ (104)
This implies that the ironed virtual value is given by (102), where $w^{**}=\hat{G}_{K}^{-1}(q^{**})$.
@@
$\frac{\displaystyle\int_{-\infty}^{q}\vartheta(\hat{G}_{K}^{-1}(r);\hat{G}_{K})dr}{q}=\vartheta(\hat{G}_{K}^{-1}(q)).$ (105)
@@
By Theorem 2.8 of Hartline and Roughgarden (2008), an efficient mechanism maximizes the value of $\hat{x}(w)$ in regions where $\bar{\vartheta}(w;\hat{G}_{K})$ is larger while fulfilling the resource constraint and the unit demand condition. The conclusion is immediate from this. ∎
@@ B.8 Proof of Theorem 8
By Proposition 3, for any $0<\underline{c}<\bar{c}<\infty$, $\hat{G}_{K}\to\Phi_{\alpha}$, $\hat{g}_{K}\to\phi_{\alpha}$, $\hat{g}^{\prime}_{K}\to\phi^{\prime}_{\alpha}$ uniformly on $\underline{c},\bar{c}]$. Since the hazard-rate derivative $r^{\prime}(w;F)$ is a continuous functional of $(F,f,f^{\prime})$ as long as 1F1-F is bounded away from 0, we obtain $r^{\prime}(w;\hat{G}_{K})\to r^{\prime}(w;\Phi_{\alpha})$ uniformly on $\underline{c},\bar{c}]$.
Since $\Phi_{\alpha}$ has an IDHR, there exists $w^{*}>0$ such that $r^{\prime}(w;\Phi_{\alpha})>0$ for $w<w^{*}$ and $r^{\prime}(w;\Phi_{\alpha})<0$ for $w>w^{*}$. We choose $\underline{c}$ and $\bar{c}$ such that $0<\underline{c}<\min\{\epsilon,w^{*}\}$ and $\bar{c}>w^{**}>w^{*}$, where $w^{**}$ is the unique solution to (24). Then, for any small $\delta>0$, for all sufficiently large $K$, $r^{\prime}(w;\hat{G}_{K})>0$ on $\underline{c},w^{*}-\delta]$ and $r^{\prime}(w;\hat{G}_{K})<0$ on $w^{*}+\delta,\bar{c}]$.
Since $\vartheta(\cdot;\hat{G}_{K})$ is decreasing on $\underline{c},w^{*}-\delta)$, the efficient allocation rule $\hat{x}$ must be constant on an interval starting from (or before) $\underline{c}$. Specifically, there exists $w_{K}^{\dagger}>w^{*}-\delta$ and a constant $\bar{x}_{K}\in[0,1]$ such that $\hat{x}(w)=\bar{x}_{K}$ for all $w\in[\underline{c},w_{K}^{\dagger})$, and $\hat{x}(w)>\bar{x}_{K}$ for $w>w_{K}^{\dagger}$.
To evaluate the location of $w_{K}^{\dagger}$, we define the auxiliary function $\Theta$. For a distribution $\hat{G}$, let
@@
Note that $w^{**}$ is the unique solution to $\Theta(w;\Phi_{\alpha})=0$. Furthermore, for a lower bound $\underline{c}>0$, define the truncated version $\Theta_{\underline{c}}$ by
@@
Due to the uniform convergence of $\Theta_{\underline{c}}(\cdot;\hat{G}_{K})$ and its derivative to those of $\Theta_{\underline{c}}(\cdot;\Phi_{\alpha})$ on $(\underline{c},\bar{c}]$, combined with the IDHR property of $\Phi_{\alpha}$, $\Theta_{\underline{c}}(w;\hat{G}_{K})=0$ has a unique solution $w_{K,\underline{c}}^{\dagger\dagger}$ for sufficiently large $K$. Analogous to the proof of Lemma 3, the definition of $w_{K,\underline{c}}^{\dagger\dagger}$ ensures that the line segment connecting $(\hat{G}_{K}(\underline{c}),H(\hat{G}_{K}(\underline{c})))$ and $(\hat{G}_{K}(w_{K,\underline{c}}^{\dagger\dagger}),H(\hat{G}_{K}(w_{K,\underline{c}}^{\dagger\dagger})))$ is tangent to $H$ at $w_{K,\underline{c}}^{\dagger\dagger}$. Due to the IDHR structure, this line segment lies below $H$ on the interval $\underline{c},w_{K,\underline{c}}^{\dagger\dagger}]$. Since this line segment connects two points on the graph of $H$, the ironed cumulative virtual value $I$, being the global convex hull of $H$, must lie below it. Consequently, $I$ is strictly less than $H$ on the interior of the interval. By the property of the convex hull, $I$ must be a line segment on any interval where it strictly detaches from $H$. Thus, the ironed virtual value is constant on $\underline{c},w_{K,\underline{c}}^{\dagger\dagger}]$. Since $w_{K}^{\dagger}$ is the right endpoint of the constant allocation interval starting from (or before) $\underline{c}$ for the original distribution, it must satisfy $w_{K}^{\dagger}\geq w_{K,\underline{c}}^{\dagger\dagger}$.
We now show that $w_{K,\underline{c}}^{\dagger\dagger}$ approaches $w^{**}$ for sufficiently small $\underline{c}$ and large $K$. Since $\Phi_{\alpha}$ has an IDHR, $\Theta(\cdot;\Phi_{\alpha})$ is strictly decreasing in a neighborhood of $w^{**}$. Thus, for the fixed $\epsilon$, there exists $\eta>0$ such that
$\Theta(w^{**}-\epsilon/2;\Phi_{\alpha})>\eta\quad\text{and}\quad\Theta(w^{**}+\epsilon/2;\Phi_{\alpha})<-\eta.$ (108)
First, consider the approximation by $\Theta_{\underline{c}}$. As $\underline{c}\to 0$, $\int_{\underline{c}}^{w}(1-\Phi_{\alpha}(t))dt\to\int_{0}^{w}(1-\Phi_{\alpha}(t))dt$ and $\Phi_{\alpha}(\underline{c})\to 0$. For $w$ in the neighborhood of $w^{**}$, $\Phi_{\alpha}(w)-\Phi_{\alpha}(\underline{c})$ is strictly positive and bounded away from zero for sufficiently small $\underline{c}$. Thus, $\Theta_{\underline{c}}(w;\Phi_{\alpha})\to\Theta(w;\Phi_{\alpha})$ for any $w$ in the neighborhood of $w^{**}$. We can choose $\underline{c}<\min\{\epsilon,w^{*}\}$ small enough such that for $w\in\{w^{**}-\epsilon/2,w^{**}+\epsilon/2\}$,
$|\Theta_{\underline{c}}(w;\Phi_{\alpha})-\Theta(w;\Phi_{\alpha})|<\eta/3.$ (109)
Fix such $\underline{c}$.
Next, consider the convergence as $K\to\infty$. Since $\hat{G}_{K}\to\Phi_{\alpha}$ uniformly on $\underline{c},\bar{c}]$ and $\Phi_{\alpha}(w)-\Phi_{\alpha}(\underline{c})$ is bounded away from zero for $w$ in the neighborhood of $w^{**}$, $\Theta_{\underline{c}}(\cdot;\hat{G}_{K})$ converges uniformly to $\Theta_{\underline{c}}(\cdot;\Phi_{\alpha})$ on the neighborhood of $w^{**}$. Consequently, there exists $K_{0}$ such that for all $K>K_{0}$,
$|\Theta_{\underline{c}}(w;\hat{G}_{K})-\Theta_{\underline{c}}(w;\Phi_{\alpha})|<\eta/3\quad\text{for }w\in\{w^{**}-\epsilon/2,w^{**}+\epsilon/2\}.$ (110)
Combining these evaluations, for $K>K_{0}$:
$\Theta_{\underline{c}}(w^{**}-\epsilon/2;\hat{G}_{K})$ $>\Theta(w^{**}-\epsilon/2;\Phi_{\alpha})-2\eta/3>\eta/3>0,$ (111)
$\Theta_{\underline{c}}(w^{**}+\epsilon/2;\hat{G}_{K})$ $<\Theta(w^{**}+\epsilon/2;\Phi_{\alpha})+2\eta/3<-\eta/3<0.$ (112)
By the intermediate value theorem, the root $w_{K,\underline{c}}^{\dagger\dagger}$ lies in $(w^{**}-\epsilon/2,w^{**}+\epsilon/2)$.
Finally, we have
$w_{K}^{\dagger}\geq w_{K,\underline{c}}^{\dagger\dagger}>w^{**}-\epsilon/2>w^{**}-\epsilon.$ (113)
Since $\hat{x}(w)=\bar{x}_{K}$ for $w\in[\underline{c},w_{K}^{\dagger})$ and $\underline{c}<\epsilon$, it follows that $\hat{x}(w)=\bar{x}_{K}$ for all $w\in(\epsilon,w^{**}-\epsilon)$. ∎